# The Coase Theorem and Ideal Exchanges

### v4

@@ 4. Secondary Results
Due to formal necessity, throughout the following we shall work within a fixed economy $\mathcal{E}$ on which every $\mathcal{I}$ operates, where $\mathcal{E}$ is defined as the 3-tuple $(\mathbb{A},R,W)$.
@@
Suppose, to the contrary, that there exists some initial distribution $\big{(}\mathcal{I}_{0}(a_{i})\big{)}_{i=0}^{n-1}\in\mathcal{P}(R)^{n}$ such that for all $t\in\mathbb{N}$, there is a $k\in\mathbb{N}$ where $\big{(}\mathcal{I}_{t}(a_{i})\big{)}_{i=0}^{n-1}\neq\big{(}\mathcal{I}_{t+k}(a_{i})\big{)}_{i=0}^{n-1}$. In other words, every “stable” distribution that follows from this initial distribution is temporary. There cannot be a case where a distribution is constant for at least two instances before undergoing a non-trivial change; suppose, to the contrary, that such a phenomenon occurred, then the first distribution at $t$ is either one where no agent is willing to trade with another agent, or where no agent is able to trade with another agent. If a break in stagnation occurs at $t+k$, then both conditions must be satisfied at $t+k-1$, but this is contradictory because the distribution at $t+k-1$ descended from $t$. It follows that $\forall t\in\mathbb{N}\Big{[}\big{(}\mathcal{I}_{t}(a_{i})\big{)}_{i=0}^{n-1}\neq\big{(}\mathcal{I}_{t+1}(a_{i})\big{)}_{i=0}^{n-1}\Big{]}$. By (B1), no exchange can be reversed. Therefore, for indefinitely distinct trades, either the bundles are always different or the agents are always different. Since there are only finitely many resources ($m$) and finitely many agents ($n$), there are finitely many possible distributions ($n^{m}$) that dually comply with (A1) and (A2); hence, the existence of at least one cycle is guaranteed when $t=n^{m}$. [^6] This is due to the pigeonhole principle. By (B2), every cycle is reducible to a reversal of one’s preferences, thereby contradicting asymmetry. It follows that such an initial distribution cannot exist. ∎
@@
Firstly, observe that (B1), (B2) and (B3) characterize a linear ordering of all elements in $\mathcal{P}(R)$. For each individual agent $a_{q}$, we reduce these preferences to a $2^{m}$-tuple through $\mathfrak{P}:\mathbb{A}\to\mathcal{P}(R)^{|\mathcal{P}(R)|}$, where $\mathfrak{P}(a_{q})=\big{(}\alpha_{i}\big{)}_{i=0}^{2^{m}-1}$, such that the following holds:
@@
We may now construct an explicit counterexample to 4.5. Suppose that $\mathbb{A}=\{a_{0},a_{1},a_{2}\}$, $R=\{x,y,z\}$, and the following statements are satisfied:
@@ 5. Primary Result
Suppose we accept every assumption in Section 3 aside from (C1) and (C2). Instead, we shall adopt the following in their place for all $t\in\mathbb{N}$:
@@
We may now prove the Coase theorem.
@@ 6. Philosophical & Concluding Remarks
Firstly, observe that Propositions 4.3 and 4.5 are formal variants of how the Coase theorem can be informally interpreted. The contention that every initial distribution converges to one unique outcome is a particularly strong claim; it is generally false because if every person endeavors to maximize what they own, and for each initial distribution there is one distinct agent who is in possession of all resources, then the ultimate outcome is both Pareto optimal and halts on this initial input irrespective of what this input may be. The invariance of the ultimate distribution is a guaranteed phenomenon in extremely narrow circumstances where there are only two agents, a single resource, and their preferences are each other’s mirror image (i.e., if $a_{0}$ prefers $x$ over its absence, then $a_{1}$ necessarily prefers its absence over $x$.)
We now turn to the primary disproof in 4.6 where we offer an explicit counterexample to Proposition 4.5. Evidently, it would seem that the exchange mechanism prescribed by (C1) is too strong, and it has been argued that the inflexibility of a double coincidence of wants is a problem in markets where exchanges are conducted through bartering. The problem is certainly not unique to such societies and has not fully disappeared in markets where there exists a dominant medium of exchange (i.e., where it has only morphed into the need for a double coincidence of value.) To see this, we may turn to 4.6 and reinterpret $z$ as a $20\$$ bill, whereas $x$ and $y$ continue to stand for material goods. $a_{0}$ still prefers $y$ over $x$ and $x$ over $20\$$: $a_{1}$ prefers $x$ over $20\$$ and $20\$$ over $y$, while $a_{2}$ prefers $20\$$ over $y$ and $y$ over $x$. If the initial allocation were ($\big{(}\{x\},\{20\$\},\{y\}\big{)}$, then $a_{1}$ cannot purchase $x$ from $a_{0}$, who values it over $20\$$. It follows from 4.6 that no exchange occurs and the distribution is suboptimal. This is a mathematically trivial demonstration, yet its philosophical significance for economic theory lies in the observation that a market exchange should not be viewed as a private activity among only two parties. If it were, then it would be irrational for $a_{1}$ to purchase $y$ from $a_{2}$ with $20\$$, as that would make her worse off immediately. But if she chooses to do so and subsequently barters $y$ for $x$ with $a_{0}$, then everyone is made better off. Naturally, it would be more efficient if all three agents recognized this, and immediately arrived at a better distribution through a 3-party agreement. The presence of a common medium of exchange has a non-trivial effect if we assume that preferences are incomplete over resources but complete over the medium of exchange; we reject this two-sorted-ness as it is irrational and would contradict (B3). Furthermore, it is inconsistent with (B2) because the defining quality of a medium of exchange is its ability to measure the value of an item. If we can compare an item to a quantity of the medium, and compare that quantity with the quantity we derive from some other comparable item, then we should be able to compare the items themselves as though we lived in a bartering economy. [^7] Let $R$ be a set of resources such that it consists of a unique multiset of identical goods $\mathcal{N}\subseteq R$ for which every agent has quasi-linear preferences: $\forall a_{q}\in\mathbb{A}\forall A,B\in\mathcal{P}(R)\big{[}(((A\setminus(A\cap B)\subseteq\mathcal{N})\land(B\setminus(A\cap B)\subseteq\mathcal{N}))\land(|A\setminus(A\cap B)|<|B\setminus(A\cap B)|))\implies(A,B)\in W(a_{q})\big{]}.$ The existence of this numeraire does not solve the predicament since agents may freely elect to not have a price tag on particular material resources. In other words, no amount of money would be enough to compel them to sell off what they own (i.e., $\Diamond\,\exists a_{q}\in\mathbb{A}\exists A\in\mathcal{P}(R)\forall B\subseteq\mathcal{N}[(B,A)\in W(a_{q})]$). If one were to contend that a price tag can be assigned to every bundle of resources in $\mathcal{P}(R)$ (i.e., $\Box\,\forall a_{q}\in\mathbb{A}\,\forall A\in\mathcal{P}(R)\,\exists B\subseteq\mathcal{N}[(A,B)\in W(a_{q})]$) then this would call into question the value of the medium. Evidently, the numeraire’s worth is contingent on the value of the goods it can buy: it cannot be pursued as a bundle of resources in its own right, rather, there must exist some bundle of resources such that that bundle is strictly preferred over the numeraire. But this would create a vicious cycle: since there must exist a price tag for that bundle, there must also be a different corresponding bundle for that price.
Similar to a Dedekind cut, the notion of an ideal exchange (D1) is a precise solution to the predicament described in 4.6. As opposed to performing exchanges on the basis of pairwise preferences, we allow an arbitrary number of agents $(n)$ to form an $n$-party agreement iff an agent is made better off without there being another agent who is necessarily made worse off. Moreover, the distribution effectively halts once a Pareto optimal outcome is reached. From there, a proof of the Coase theorem seems almost tautological, since every ideal exchange is not merely an endeavor to reach a Pareto optimal distribution through the blueprint of agent preferences, but rather, is, in it of itself, defined in terms of Pareto optimality.
Finally, it would seem that the institutional safeguards of having well-defined ownership rights (A1) and the absence of unclaimed resources (A2) are insufficient for a proof of the Coase theorem, even if we assume that preferences are characteristically rational and transactions are frictionless. The ultimate question remains unresolved: what ought to be done given these preferences and this initial distribution? Philosophically, an ideal exchange can be characterized as a situation where every agent has sufficient entrepreneurial zeal or intellectual foresight to recognize and exploit the inefficiency of their current situation. This is especially evident for $a_{1}$ when she makes an exchange that does not reflect her authentic preferences in order to reach a more desirable, permanent outcome. It follows that the truth of the theorem rests on whether we should extend this presumption of foresight and entrepreneurial courage to any market participant at all.
@@ Acknowledgments
The author is grateful to Noah Wang (University of St Andrews) and Professor Benedict Eastaugh (University of Warwick) for their lasting inspiration and guidance.

### v6

@@ 4. Secondary Results
Given the model above, we can now ask: will the agents necessarily reach a Pareto optimal allocation? Is the final allocation unique? How does it depend on the initial allocation? In this section, we derive some general results about the trading process under assumptions (A1)(C2). We show that the process always converges to a stable allocation in finitely many steps. However, two commonly expected “Coasean” properties do not hold in this basic setting: (1) the final outcome can depend on the initial allocation, and (2) the final outcome need not be Pareto optimal. These results will motivate the refinement of our model in the next section. We begin by working within a fixed economy $\mathcal{E}$ on which every $\mathcal{I}$ operates, where $\mathcal{E}$ is defined as the 3-tuple $(\mathbb{A},R,W)$.
@@
Suppose, to the contrary, that there exists some initial distribution $\big(\mathcal{I}_{0}(a_{i})\big)_{i=0}^{n-1}\in\mathcal{P}(R)^{n}$ such that for all $t\in\mathbb{N}$, there is a $k\in\mathbb{N}$ where $\big(\mathcal{I}_{t}(a_{i})\big)_{i=0}^{n-1}\neq\big(\mathcal{I}_{t+k}(a_{i})\big)_{i=0}^{n-1}$. In other words, every “stable” distribution that follows from this initial distribution is temporary. There cannot be a case where a distribution is constant for at least two instances before undergoing a non-trivial change; suppose, to the contrary, that such a phenomenon occurred, then the first distribution at $t$ is either one where no agent is willing to trade with another agent, or where no agent is able to trade with another agent. If a break in stagnation occurs at $t+k$, then both conditions must be satisfied at t+k1t+k-1, but this is contradictory because the distribution at t+k1t+k-1 descended from $t$. It follows that $\forall t\in\mathbb{N}\Big[\big(\mathcal{I}_{t}(a_{i})\big)_{i=0}^{n-1}\neq\big(\mathcal{I}_{t+1}(a_{i})\big)_{i=0}^{n-1}\Big]$. By (B1), no exchange can be reversed. Therefore, for indefinitely distinct trades, either the bundles are always different or the agents are always different. Since there are only finitely many resources ($m$) and finitely many agents ($n$), there are finitely many possible distributions ($n^{m}$) that dually comply with (A1) and (A2); hence, the existence of at least one cycle is guaranteed when $t=n^{m}$. [^8] This is due to the pigeonhole principle. By (B2), every cycle is reducible to a reversal of one’s preferences, thereby contradicting asymmetry. It follows that such an initial distribution cannot exist. ∎
@@
Remark 4.5.
This simple example shows that, without additional assumptions, such as the ability to make side payments or quasi-linear utilities to ensure transferable utility, the Coasean outcome need not be unique. The initial property rights allocation can matter for who ends up with the good, even though in both cases the outcome is Pareto optimal. The example may seem trivial, but it underscores that Coase’s theorem in its strongest form – that the same efficient outcome will occur regardless of how rights are initially assigned – is not guaranteed in our model except under special conditions. Indeed, Hurwicz’s result mentioned earlier formalizes exactly such conditions.
Next, we examine whether the process described will necessarily reach a Pareto optimal allocation. One might assume that because every trade we allow is Pareto-improving, the process of making such trades until none are left must yield a Pareto-optimal outcome. However, the following proposition and theorem show that this is not always the case: the process can get stuck in a stable allocation that is not Pareto optimal, due to the bilateral nature of trades we assumed. In other words, there may exist allocations where no pair of agents can improve through trade, yet a group of three or more of agents could jointly redistribute resources in a way that benefits at least one of them without hurting anyone. Such an allocation would be a stable endpoint of bilateral trades but not Pareto efficient.
@@
Firstly, observe that (B1)(B3) characterize a linear ordering of all elements in $\mathcal{P}(R)$. For each individual agent $a_{q}$, we reduce these preferences to a $2^{m}$-tuple through $\mathfrak{P}:\mathbb{A}\to\mathcal{P}(R)^{|\mathcal{P}(R)|}$, where $\mathfrak{P}(a_{q})=\big(\alpha_{i}\big)_{i=0}^{2^{m}-1}$, such that the following holds:
@@
We may now construct an explicit counterexample to 4.6. Suppose that $\mathbb{A}=\{a_{0},a_{1},a_{2}\}$, $R=\{x,y,z\}$, and the following statements are satisfied:
@@
Remark 4.8.
The counterexample highlights another key insight: the structure of the bargaining process matters crucially for efficiency. In that example, there was nothing in our rules that allowed $a_{0}$, $a_{1}$, and $a_{2}$ to recognize and execute the three-way swap $x\to a_{1},\;y\to a_{0},\;z\to a_{2}$} all at once. The Coase theorem’s informal logic assumes that any and all opportunities for mutually beneficial rearrangements will be exploited, but if agents can only trade bilaterally, some opportunities which involve larger coalitions might not be accessible. This suggests that if we allow sufficiently many agents to cooperate simultaneously, we might restore efficiency.
@@ 5. Primary Result
The failure of bilateral bargaining to guarantee Pareto optimality leads us to refine our model. We introduce the concept of an ideal exchange, which is a trade possibly involving multiple agents and resources, chosen specifically because it makes at least one agent better off while making no one worse off. In essence, instead of requiring trades to be strictly beneficial to each participant in isolation (as in (C1)), we will now allow a trade if it benefits someone and harms no one. Such an exchange could, for example, involve three parties transferring goods in a cycle, as in the previous counterexample – something no two of them would do alone, but which all three together can agree on because it benefits each of them or at least leaves them no worse. One might think of this as agents being farsighted or capable of forming coalitions to achieve a Pareto improvement, even if any subset of that coalition wouldn’t move unilaterally.
Formally, we modify the trading rule (C1) to a new rule (D1) that allows any coalition of agents to execute an exchange that results in a Pareto better distribution. We also adjust (C2) to (D2) to define stagnation when no Pareto-improving exchange exists.
@@
Remark 5.1.
These rules (D1)(D2) capture the idealized assumption that the agents can and will reorganize any time it is possible to make someone better off without making anyone worse off, as if any coalition of agents that could benefit at no one’s expense instantly forms and executes the required deal. This might sound unrealistic in practice – it assumes costless, fully informed multilateral bargaining – but it represents the theoretical limit of what frictionless negotiation could achieve.
We now state and prove the main theorem under the revised rules, which essentially says: If property rights are well-defined, there are no transaction costs, and if agents are able to execute any exchange that is beneficial to at least one and harmful to none, then the allocation of resources will converge to a Pareto optimal state and remain there. This is a formal version of the Coase theorem.
@@
It might appear that by building Pareto optimality into the very rule (D1), we have proven the Coase theorem almost by definition – indeed, we have essentially assumed that agents have the capability to directly implement any Pareto improvement. This assumption is admittedly strong, but it isolates the core conceptual requirement for Coase’s vision to hold: agents must be able to coordinate exchanges beyond simple pairwise trades whenever such coordination can make everyone better off. In reality, achieving such coordination may involve bargaining institutions, intermediaries, or enforceable contracts among multiple parties. However, Coase’s theorem is an ideal benchmark – it presumes those coordination issues away by assuming zero transaction costs and perfectly defined rights. Our formalism with ideal exchanges is an accurate mirror of that presumption.
@@ 6. Philosophical & Concluding Remarks
What, then, is the philosophical takeaway? One observation is that Coase’s theorem is not a trivial truth of logic or economics; it holds only under specific conditions about the process of interaction. By formalizing the theorem, we made explicit some often implicit assumptions. The notion of “well-defined property rights” and “no transaction costs” were modeled by saying every resource is always owned and, initially in Section 4, by limiting trades to those both parties strictly benefit from. These were not enough to guarantee efficiency. We had to strengthen “no transaction costs” to an extreme degree – allowing costless multilateral bargaining – to get the efficiency result. This reveals that the Coase theorem’s truth hinges on assuming away not just pricing or negotiation costs, but also the complexities of coalition formation.
Our formal model also provides a new perspective on the role of entrepreneurial foresight and coordination. In the philosophical literature on economics, much is made of the idea of agents as optimization machines who will find any available improvement. Coase’s view of bargaining could be seen as assuming agents have the farsightedness and perhaps the entrepreneurial spirit to realize complex exchanges. In our terms, an ideal exchange requires that agents are aware of the collective benefit and can trust that no one will be harmed by participating. One might say: if agents truly recognize a three-way trade is to all of their benefit, they have a strong incentive to implement it, perhaps through a sequence of trades or a multilateral contract. Thus, whether we extend this presumption of foresight and entrepreneurial courage to any market participants is ultimately an empirical or at least a pragmatic question. Coase’s theorem does not hold come what may; it holds if people can bargain in an idealized way.
Finally, stepping back to the broader motivation: we used set theory and logic to axiomatize a slice of economic theory. Was this effort worthwhile? We believe it was, for several reasons. It forced us to clarify subtle concepts like what exactly constitutes a “transaction cost” in a model. Here, anything that prevents a Pareto-improving move, including the inability of three people to bargain together, can be interpreted as a transaction cost or impediment. It also allowed us to uncover assumptions that might otherwise be overlooked – for example, the standard textbook treatment with an Edgeworth box implicitly assumes that the contract curve will be reached, but fails to spell out an effectively calculable procedure.
By constructing one and seeing it fail, we identified that a gap exists for multi-party exchanges. This is a case where the philosophy of mathematics – in the sense of scrutinizing the foundations and logical structure of a theory – contributes to economic understanding. We have not proven anything that could not be intuited (indeed, our formal results echo known insights from economics), but we have provided a formal verification and clarity. Recalling Cohen, the aim is not to discover new truths per se, but to precisely articulate the rules and assumptions in use. In our case, we articulated the “rules of the game” that lead to Coasean outcomes or their absence.
In conclusion, the Coase theorem can be affirmed as a logically valid result if we assume ideal conditions for bargaining. The presence of transaction costs, broadly interpreted to include any impediment to comprehensive agreements, can invalidate the theorem. None of this undermines Coase’s fundamental insight that when people can bargain freely, they will tend to remove inefficiencies – but it does remind us that “bargaining freely” is doing a lot of work in that sentence. The formalization presented helps pinpoint what exactly “freely” must entail. It is our hope that such formal exercises, while abstract, can inform practical discussions by illuminating why certain real-world situations deviate from the ideal and what institutional arrangements might be needed to approach the efficient outcomes that the Coase theorem predicts under its stringent conditions.
@@ Acknowledgments
The author is grateful to Noah Wang (University of St Andrews) and Professor Benedict Eastaugh (University of Warwick) for their lasting inspiration and guidance. The author also thanks Professor Joel David Hamkins (University of Notre Dame) for his helpful comments.
Declarations
The author declares no conflicts of interest. No funding was received for this work. This work is purely theoretical and does not involve the generation of any data sets.